OpenStudy (mendicant_bias):

I'm not sure what I'm doing wrong here, just a simple error somewhere, but I'm trying to show by partial fraction decomposition that tanh^-1(x) = 1/2ln((1+x)/(1-x)), but I'm running into some trouble and somehow the arguments for my natural logarithms are the negative of what they should be. My steps posted below in a few minutes.

OpenStudy (mendicant_bias):

\[\frac{ d }{ dx }\tanh ^{-1}(x) = \frac{ 1 }{ 1-x ^{2} };\] \[\frac{ 1 }{ 1-x ^{2} }=\frac{ A }{ (1+x) }+\frac{ B }{ (1-x) }.\] \[\frac{ 1(1+x)(1-x) }{ (1+x)(1-x) } = \frac{ A(1+x)(1-x) }{ (1+x) }+\frac{ B(1-x)(1+x) }{ (1-x) }\]

OpenStudy (mendicant_bias):

\[1 = A(1-x)+B(1+x)\] Since this is an identity, all values of x are valid; I'll assume x = 1. \[1 = A(1-1)+B(1+1);\]\[1 = 2B; \]\[B = \frac{ 1 }{ 2 }\] Doing the same for x where x = -1 results in A = -1/2. (cont'd)

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\frac{ (-1/2) }{ (1-x) }+ \int\limits_{}^{}\frac{ (1/2) }{ (1+x) } = -\frac{ 1 }{ 2 }\int\limits_{}^{}\frac{ 1 }{ 1-x }+\frac{ 1 }{ 2 }\int\limits_{}^{}\frac{ 1 }{ 1+x }=\]

OpenStudy (mendicant_bias):

\[-\frac{ 1 }{ 2 }\ln(1-x)+\frac{ 1 }{ 2 }\ln(1+x) = \frac{ 1 }{ 2 }\ln \frac{ (1+x) }{ (1-x) }\] Wait, what? I think I just did it right, rofl.

OpenStudy (mendicant_bias):

Nope, I didn't, I switched the denominators of A and B.

OpenStudy (mendicant_bias):

But somewhere in there, if the denominators are correct, I'm ending up with the wrong result where the argument of the log is the reciprocal of what it should be.

OpenStudy (mendicant_bias):

\[-\frac{ 1 }{ 2 }\ln(1+x)+\frac{ 1 }{ 2 }\ln(1-x) = \frac{ 1 }{ 2 }\ln \frac{ (1-x) }{ (1+x) }\]????

hartnn (hartnn):

what is the derivative of inverse tan hyperbolic function ?

OpenStudy (mendicant_bias):

At the top, 1/(1-x^2).

hartnn (hartnn):

oh, k ... and the condition for that is |x| <1

OpenStudy (mendicant_bias):

Yeah, on that condition. Unless I'm losing my marbles,

OpenStudy (mendicant_bias):

I'm messing up hardcore somewhere, I just can't see where

hartnn (hartnn):

during integral of 1/1-x you should also divide by the co-efficient of x, that is -1

OpenStudy (mendicant_bias):

I don't know what you mean when you mentioned dividing by the coefficient of x during the integral, what?

hartnn (hartnn):

int (1/2)/(1-x) dx = 1/2 (ln (1-x)) * (-1)

hartnn (hartnn):

like \(\int \dfrac{1}{ax+b}dx =(1/a)\ln |ax+b|+c\)

hartnn (hartnn):

because d/dx (ln[ax+b]) = a/ax+b

OpenStudy (mendicant_bias):

Oh! Derp. Okay, wait just one second.

hartnn (hartnn):

and A comes out to be +1/2 -_-

hartnn (hartnn):

1= A (1- (-1)) 1=A(1+1)

OpenStudy (mendicant_bias):

I dun goof'd.

OpenStudy (mendicant_bias):

\[x = 1\] \[1 = A(1-x) + B(1+x); \]\[1 = A(1-(1))+B(1+(1));\]\[1 = 2B; \]\[B = \frac{ 1 }{ 2 }\] \[x = (-1)\] \[1 = A(1-(-1)) + B(1+(-1));\]\[1 = 2A;\]\[A = \frac{ 1 }{ 2 }.\]

OpenStudy (mendicant_bias):

(Cont'd for logarithm/integration part)

hartnn (hartnn):

its correct, now u will get it

OpenStudy (mendicant_bias):

\[\int\limits_{}^{}\frac{ (1/2) }{ (1+x) }dx+\int\limits_{}^{}\frac{ (1/2) }{ (1-x) }dx = \frac{ 1 }{ 2 }\int\limits_{}^{}\frac{ 1 }{ 1+x }+\frac{ 1 }{ 2 }\int\limits_{}^{}\frac{ 1 }{ 1-x }\]\[\frac{ 1 }{ 2 }\ln(1+x)+\frac{ 1 }{ 2 }(-1)\ln(1-x) = \frac{ 1 }{ 2 }\ln \frac{ (1+x) }{ (1-x) }\]

OpenStudy (mendicant_bias):

Huzzah.

hartnn (hartnn):

\(\huge \checkmark\)

OpenStudy (mendicant_bias):

How did you do that? (check)

OpenStudy (mendicant_bias):

I know you can use LaTeX to write equations here, was it that, or something else?

hartnn (hartnn):

latex only `\(\checkmark\)`

hartnn (hartnn):

try to give it a color too

hartnn (hartnn):

\(\huge \color{red} \checkmark\) \(\huge \color{green} \checkmark\)